# Empirical Analysis of the Model-Free Valuation Approach: Hedging Gaps, Conservatism, and Trading Opportunities

### v1

@@ 5.2. Testing the strategy
Since we do not have access to historical prices of OTC-traded exotic derivatives, as a realistic surrogate we simulate derivative prices $\mathfrak{M}(\Phi)$ by a calibrated Heston-model as described in Appendix D. Due to the model-free construction, using derivative prices from a Heston model instead of market quotes only changes the number of trades being triggered but not the average profit conditional on trades being initiated.
@@
Figure 8. Left: The graphs show $\underline{q}(c)$ and the the average returns of $\frac{\Phi(S)-\underline{u_{d,(\theta_{i,j}),(H_{i})}}(S)-\underline{C}}{S_{0}}$ conditional on $\frac{\mathfrak{M}(\Phi)-\underline{\operatorname{P}}}{S_{0}}\leq c$ in dependence of $c$, cost $\underline{C}=\underline{\operatorname{P}}-\mathfrak{M}(\Phi)$, evaluated on training data.
@@
Right: The graphs show $\overline{q}(c)$ and the the average returns of $\frac{\overline{u_{d,(\theta_{i,j}),(H_{i})}}(S)-\Phi(S)-\overline{C}}{S_{0}}$ conditional on $\frac{\overline{\operatorname{P}}-\mathfrak{M}(\Phi)}{S_{0}}\leq c$ in dependence of $c$, cost $\overline{C}=\mathfrak{M}(\Phi)-\overline{\operatorname{P}}$, evaluated on training data.
@@
Figure 8 illustrates the dependence of $\underline{q}(c)$ and $\overline{q}(c)$ on $c$ as well as the average returns of $\frac{\Phi(S)-\underline{u_{d,(\theta_{i,j}),(H_{i})}}(S)-\underline{C}}{S_{0}}$ conditional on ${\mathfrak{M}(\Phi)-\underline{\operatorname{P}}(\Phi)}\leq c\cdot S_{0}$, i.e., conditional on a trade being triggered as the market price is close enough to the lower model-free price bound, where we denote the costs of this strategy by $\underline{C}:=\mathfrak{M}(\Phi)-\underline{\operatorname{P}}(\Phi)$. Analogue, we study $\frac{\overline{u_{d,(\theta_{i,j}),(H_{i})}}(S)-\Phi(S)-\overline{C}}{S_{0}}$ conditional on ${\overline{\operatorname{P}}(\Phi)-\mathfrak{M}(\Phi)}\leq c\cdot S_{0}$ corresponding to the situation that a trade is triggered since the market price is close enough to the upper bound, where $\overline{C}:=\overline{\operatorname{P}}(\Phi)-\mathfrak{M}(\Phi)$. We evaluate on the training data. The two subplots at the bottom of Figure 8 show how the average net profit varies when gradually increasing $c$. Note that in this consideration the costs of the strategy are taken into account which explains the observation that with an increasing $c$ the average profit decreases due to the increasing costs of the strategy.
@@
c -0.0047 -0.0001 0.0014 0.0060 0.0123 0.0200 0.0300
Number of Trades 637 1142 1420 2653 4376 6502 9211
Avg return p.a. 2855.59 2350.08 2038.54 1293.50 858.34 632.62 447.80
Std dev. p.a. 367.10 593.99 536.26 408.17 326.35 327.66 276.96
Down dev. p.a. 7.43 31.14 15.88 20.47 24.66 26.36 26.90
Sharpe ratio 7.78 3.96 3.80 3.17 2.63 1.93 1.62
Sortino ratio 384.35 75.47 128.34 63.20 34.80 24.00 16.64
Max. Loss 21.11 22.58 22.58 24.51 52.14 67.00 143.69
Max. Profit 1073.39 3289.63 3289.63 3289.63 3289.63 13854.70 13854.70
@@
c -0.0147 -0.0082 -0.0049 0.0009 0.0012 0.0049 0.0108
Number of Trades 0 0 0 2 2 4 11
Avg return p.a. -15.14 -15.14 -27.83 129.35
Std dev. p.a. 1.80 1.80 1.39 28.64
Down dev. p.a. 0.19 1.25
Sharpe ratio -8.42 -8.42 -20.02 4.52
Sortino ratio -146.13 103.86
Max. Loss 2.94 2.94 20.46 20.46
Max. Profit 9.97 9.97 9.97 374.60
@@
$\overline{q}(c)$ 0.95 0.85 0.62 0.31 0.09
$\underline{q}(c)$ 1.00 0.99 0.94 0.15 0.14
Number of Trades 1346 4034 11848 41972 48211
Avg return p.a. 3059.80 133.84 667.53 327.86 336.49
Std dev. p.a. 577.92 382.87 310.50 304.07 530.71
Down dev. p.a. 4.78 35.49 51.42 77.43 72.23
Sharpe ratio 5.29 3.48 2.15 1.08 0.63
Sortino ratio 640.15 37.58 12.98 4.23 4.66
Max. Loss 28.44 62.68 111.82 187.33 305.37
Max. Profit 1675.88 1675.88 4120.82 15308.64 160994.61
@@ Appendix A Numerical Algorithm
     Define $\mathbb{G}_{k+1}:=\mathbb{G}_{k}\cup\{S\in\mathbb{G}\leavevmode\nobreak\ |\leavevmode\nobreak\ u_{x^{(k)}}(S)-\Phi(S)<\delta+\frac{1}{k+1}\};$
@@ Appendix D Benchmark Models
for $r$ being the interest rate, $\kappa$ the volatility mean-reversion speed, $\theta$ the volatility long term mean, $\sigma$ the diffusion parameter associated to the volatility, and where $W_{t}^{1}$ and $W_{t}^{2}$ are two correlated Brownian motions. The parameters of the Heston-model are calibrated to the considered call option prices according to the methodology described in levendorskiui2012efficient. The price of the forward start option (and based on this its sensitivities) can be computed as described in Kruse2005.

### v2

@@ 5.2. Testing the strategy
Since we do not have access to historical prices of OTC-traded exotic derivatives, as a realistic surrogate we simulate derivative prices $\mathfrak{M}(\Phi)$ by a calibrated Heston-model. We choose the Heston model as a representative stochastic-volatility benchmark, as forward-start options are particularly sensitive to forward variance dynamics; similar conclusions are expected for other well-calibrated smile-consistent models. The Heston model parameters are calibrated on a daily basis for each underlying stock using all available call option quotes, the underlying spot price, and the corresponding interest-rate inputs observed on that day (see the data description in Section 2.1 and interest rate calculation in Section 3.3.2). We estimate the parameters by minimizing the squared pricing error between market option prices (mid-quotes) and model-implied prices, and we add a Tikhonov regularization term to stabilize the parameter estimates over time. We implement the calibration in Python using QuantLib (quantlib), which provides an analytic European-option Heston pricer (semi-closed-form/Fourier-based). Further details of the calibration are described in Appendix D. Then, using the calibrated parameters, we simulate daily market prices via Monte Carlo simulation, performing 50,000 simulation paths with 1,000 time steps per option. Our Heston calibration can occasionally produce option prices that violate the no-arbitrage bounds. This is not unexpected: the empirical option data itself may contain residual arbitrage, and the calibration is further subject to numerical and approximation errors. To avoid artificially improving the reported performance by trading on model-implied mispricings that are driven by these artifacts rather than genuine opportunities, we impose a conservative filter: whenever a Heston-model price falls outside the no-arbitrage range, we discard that option and do not allow trading in it. This filtering step is conservative and biases results against finding arbitrage trades, since model-induced arbitrage opportunities are excluded by construction.
The use of surrogate model prices at this step can be justified from a practical viewpoint, as it has been well studied that after calibration, the Heston-model reproduces the broad shape of observed implied-volatility smiles/skews reasonably well—particularly away from very short maturities, see, e.g. bakshi1997empirical, fiorentini2002estimation and gatheral2011volatility. We emphasize that using prices originating from a calibrated Heston-model does not mean that we assume that the underlying security follows the dynamics of the Heston model, and the hedge itself remains model-free; however, in the absence of historical OTC prices, the empirical evaluation of the trading rule necessarily relies on surrogate prices generated from a calibrated model.
@@
Figure 8. Left: The graphs show $\underline{q}(c)$ and the average returns of $\frac{\Phi(S)-\underline{u_{d,(\theta_{i,j}),(H_{i})}}(S)-\underline{C}}{S_{0}}$ conditional on $\frac{\mathfrak{M}(\Phi)-\underline{\operatorname{P}}}{S_{0}}\leq c$ in dependence of $c$, cost $\underline{C}=\underline{\operatorname{P}(\Phi)}-\mathfrak{M}(\Phi)$, evaluated on training data.
@@
Right: The graphs show $\overline{q}(c)$ and the the average returns of $\frac{\overline{u_{d,(\theta_{i,j}),(H_{i})}}(S)-\Phi(S)-\overline{C}}{S_{0}}$ conditional on $\frac{\overline{\operatorname{P}(\Phi)}-\mathfrak{M}(\Phi)}{S_{0}}\leq c$ in dependence of $c$, cost $\overline{C}=\mathfrak{M}(\Phi)-\overline{\operatorname{P}}(\Phi)$, evaluated on training data.
@@
Figure 8 illustrates the dependence of $\underline{q}(c)$ and $\overline{q}(c)$ on $c$ as well as the average returns of $\frac{\Phi(S)-\underline{u_{d,(\theta_{i,j}),(H_{i})}}(S)-\underline{C}}{S_{0}}$ conditional on ${\mathfrak{M}(\Phi)-\underline{\operatorname{P}}(\Phi)}\leq c\cdot S_{0}$, i.e., conditional on a trade being triggered as the market price is close enough to the lower model-free price bound, where we denote the costs of this strategy (i.e., the costs of buying the option and selling the maximal sub-hedge) by $\underline{C}:=\mathfrak{M}(\Phi)-\underline{\operatorname{P}}(\Phi)$. Analogue, we study $\frac{\overline{u_{d,(\theta_{i,j}),(H_{i})}}(S)-\Phi(S)-\overline{C}}{S_{0}}$ conditional on ${\overline{\operatorname{P}}(\Phi)-\mathfrak{M}(\Phi)}\leq c\cdot S_{0}$ corresponding to the situation that a trade is triggered since the market price is close enough to the upper bound, where $\overline{C}:=\overline{\operatorname{P}}(\Phi)-\mathfrak{M}(\Phi)$ denotes the costs of this strategy, i.e., the costs of selling the option and buying the minimal super-hedge. We evaluate on the training data. The two subplots at the bottom of Figure 8 show how the average net profit varies when gradually increasing $c$. Note that in this consideration the costs of the strategy are taken into account which explains the observation that with an increasing $c$ the average profit decreases due to the increasing costs of the strategy.
@@
c 0.0002 0.0011 0.0025 0.0061 0.0113 0.0182 0.0272
Number of Trades 590 694 887 1456 2241 3220 4493
Avg return p.a. 1378.97 1280.04 1098.92 810.59 568.78 404.07 295.95
Std dev. p.a. 135.98 128.54 117.32 99.88 87.30 76.24 66.21
Down dev. p.a. 2.79 6.07 13.87 24.98 23.07 19.13
Sharpe ratio 10.14 9.96 9.37 8.12 6.52 5.30 4.47
Sortino ratio 458.99 181.03 58.46 22.77 17.51 15.47
Max. Loss 0.05 1.50 3.16 11.26 19.52 31.99 46.08
Max. Profit 388.72 388.72 388.72 388.72 388.72 388.72 388.72
@@
c 0.0000 0.0009 0.0021 0.0053 0.0093 0.0146 0.0208
Number of Trades 1 1 1 1 2 19 317
Avg return p.a. 5.05 5.05 5.05 5.05 3.07 1.27 28.70
Std dev. p.a. 0.18 0.18 89.36
Down dev. p.a. 0.02 57.52
Sharpe ratio 17.38 7.22 0.32
Sortino ratio 72.66 0.50
Max. Loss 0.94 35.33
Max. Profit 9.97 9.97 9.97 9.97 9.97 9.97 181.30
@@
Next, we consider an alternative position-sizing rule according to which we invest at all times, and we let the investment amount depend on the distance to the nearest price bound: the closer the market price is to a bound (i.e., the smaller $c$), the larger the position we take. Specifically, we scale the investment by
(5.2) $\exp\left(-\beta\cdot\left(\frac{\min\{M(\Phi)-\underline{P}(\Phi),\overline{P}(\Phi)-M(\Phi)\}}{S_{0}}\right)\right).$
Here, the term inside the minimum is the market price’s distance to the nearest no-arbitrage bound, and the parameter $\beta>0$ controls how aggressively the investment is reduced as this distance increases: a larger $\beta$ leads to a sharper decline of the scaling factor from (5.2) (i.e., investments tend to being concentrated on observations closer to the bounds). Results in the testing period can be found in Table 5. Note that the strategy trades at all times; however, by construction the position size becomes almost negligible when market prices are far from the price bounds. The optimal Sharpe ratio is attained for $\beta\approx 50$, whereas for the Sortino ratio, in line with the results from the first investment strategy, it seems to be optimal to choose $\beta$ as large as possible (meaning we prioritize strongly to invest more for prices being close to the price bound). For $\beta=100$ we report a Sortino ratio of $7.29$.
@@
$\overline{q}(c)$ 1.00 0.89 0.67 0.34 0.06
$\underline{q}(c)$ 1.00 1.00 0.91 0.00 0.00
Number of Trades 0 1607 6326 24962 29193
Avg return p.a. 630.71 515.06 386.22 353.90
Std dev. p.a. 298.68 304.28 369.80 343.07
Down dev. p.a. 42.35 54.33 74.60 68.30
Sharpe ratio 2.11 1.69 1.04 1.03
Sortino ratio 14.89 9.48 5.18 5.18
Max. Loss 40.53 111.15 242.57 303.75
Max. Profit 1341.05 2565.04 2565.04 3040.10
table
Table 5. Test period performance with exponential investment
$\beta$ 100 50 10 5 1
Number of Trades 30162 30162 30162 30162 30162
Avg return p.a. 42.08 94.56 253.58 294.25 333.02
Std dev. p.a. 41.12 84.58 241.20 284.47 326.07
Down dev. p.a. 5.78 14.61 47.49 56.28 64.68
Sharpe ratio 1.02 1.12 1.05 1.03 1.02
Sortino ratio 7.29 6.47 5.34 5.23 5.15
Max. Loss 29.56 43.03 156.25 194.69 276.84
Max. Profit 939.04 1118.76 2080.93 2310.34 2863.32
@@ 6. Conclusion
The approach presented in this paper can be easily extended to other financial instruments such as options on variance or volatility indices, e.g., VIX options or variance swap options. They provide a natural and promising setting for further market-based empirical tests of our framework, in particular, since these instruments are actively traded and therefore allow one to assess the performance of model-free hedging strategies using directly observed market prices rather than synthetically constructed ones.
To implement the strategy the static component of the hedging strategy can be calibrated to observed option prices on volatility indices or variance swaps, while the dynamic component can be formulated with respect to the corresponding futures or index levels. This illustrates that our approach is not confined to synthetic examples but extends naturally to realistic market environments with liquidly traded derivatives. This setting therefore provides a natural bridge between robust, model-free hedging theory and empirically observable volatility markets.
@@ Appendix A Numerical Algorithm
     Define $\mathbb{G}_{k+1}:=\mathbb{G}_{k}\cup\{S\in\mathbb{G}\penalty 10000\ |\penalty 10000\ u_{x^{(k)}}(S)-\Phi(S)<\delta+\frac{1}{k+1}\};$
@@ Appendix D Benchmark Models
for $r$ being the interest rate, $\kappa$ the volatility mean-reversion speed, $\theta$ the volatility long term mean, $\sigma$ the diffusion parameter associated to the volatility, and where $W_{t}^{1}$ and $W_{t}^{2}$ are two correlated Brownian motions. The parameters of the Heston-model are calibrated to the considered call option prices according to the methodology described in levendorskiui2012efficient. On each trading day, we use data of all option prices, stock prices and interest rates of this day, and calibrate Heston parameters for each underlying stock. To calibrate the Heston model parameters, we minimize the squared distance between observed market prices and model-implied prices obtained under the Heston dynamics. To enhance numerical stability and mitigate sensitivity to small perturbations in the input data, we additionally incorporate a Tikhonov-type regularization term (see, e.g., egger2005tikhonov) that penalizes abrupt temporal changes in the parameter estimates. The resulting calibration problem is formulated as
$\min_{(\Theta_{t,j})_{j=1,\dots,m}}\{\sum_{i=1}^{n}\left\lVert\operatorname{P}(\Phi_{i},\Theta_{t})-\mathfrak{M}(\Phi_{i})\right\rVert_{2}^{2}+\alpha\sum_{j=1}^{m}\left\lVert\Theta_{t,j}-\Theta_{t-1,j}\right\rVert_{2}^{2}\},$
where $\mathfrak{M}(\Phi_{i})$, denotes the market price of the $i$-th option, and $\operatorname{P}(\Phi_{i},\Theta_{t})$ is the corresponding model price under the Heston model with parameter vector $\Theta_{t}$ at time $t$. Here, $\Theta_{t}=(\Theta_{t,j})_{j=1,\dots,m}$, collects the $m=5$ Heston parameters at time $t$ and $\alpha>0$ controls the strength of regularization. Once the model parameters have been calibrated, the price of a forward-start option, and, consequently, its sensitivities, can be computed in closed form. We refer to Kruse2005 for the explicit pricing formula.